OpenStudy (anonymous):

Zacharias has a total of 65 coins, all of which are either dimes or nickels. The total value of the coins is $5.95. Find the number of each type of coin. Zacharias has a total of 65 coins, all of which are either dimes or nickels. The total value of the coins is $5.95. Find the number of each type of coin. @Mathematics

OpenStudy (anonymous):

Convert words and sentences into equations. Assign variable names for unknown items. In this example, number of nickels and number of dimes are unknown. Call them N and D. Based on the first sentence, we know that N + D = 65. Based on the second sentence, we know that N times value of nickel + D times value of dime total 5.95. Now, you have two equations in two variables. You should be able to solve the two equations to figure out N and D.

OpenStudy (anonymous):

Oh well, there you have the answer written for you.

OpenStudy (anonymous):

dime x nickel y x+y=65 10x+5y=595 y=65-x 10x+5(65-x)=595 10+325-5x=595 5x=270 x=54 54+y=65 y=11 54 dimes; 11 nickels